OpenStudy (anonymous):

Convert y=2(x-5)(x=2) into standard form, using the foil method

OpenStudy (anonymous):

y=(x-2)(x+5), y=(x-2)(x+5) y=x^2+5x-2x-10 y=x^2+3x-10 i tried..

OpenStudy (shubhamsrg):

you missed "2" on RHS anyways, test seems fine :)

OpenStudy (shubhamsrg):

rest*

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

your welcome